A business owner makes 1,000 items a day. Each day he or she contributes eight hours to produce those items. If hired, elsewhere he or she could have earned $250 an hour. The items sell for $15 each. Production does not stop on the weekends. If the explicit costs total $150,000 for 30 days, what is the economic profit for the month?
Answer #1

ECONOMIC PROFIT=TOTAL REVENUE-TOTAL COST

=P*Q-EXPLICIT COST-IMPLICIT COST

=1000*15*30-150000-250*8*30

=450000-350000-60000

=40000